# Optimal design of unimorph-type cantilevered piezoelectric energy harvesters using level set-based topology optimization by considering manufacturability

### v1

@@ 3 Formulation of constraints for manufacturability
3.1 Constraint: the same cross-sectional shape in each domain
@@
Within the framework of topology optimization obtained via the level set method, qualitative control over complexity can be regulated by the regularization coefficient $\tau$, as defined in Eq. (25) Yamada et al. (2010). To leverage this control over complexity, we propose strategies to assign regularization coefficients that curtail the complexity in the thickness direction of the structure, thereby facilitating manufacturability via etching. We provided the regularization coefficient as a tensor $\bm{\tau}$, and modified Eq. (26) as follows:
@@
Although the method described above controls the complexity of the structure throughout the design domain, the substrate and piezoelectric components can be readily processed via etching, even if their structures differ. It is desirable to control the complexity in the thickness direction within the substrate and piezoelectric domains without affecting the structural differences at the interface of each domain. To this end, in this study, we defined separate level set functions for the substrate and piezoelectric structures, invoking the extended level set method Noda et al. (2022), which is capable of handling multiple materials and updating each level set function concurrently. The fixed design domains for the substrate and piezoelectric material are denoted as $D_{sb}$ and $D_{pe}$, respectively, assuming that these domains remain invariant throughout the optimization procedure. The fixed design domain, denoted as $D$, encompasses both $D_{pe}$ and $D_{sb}$. Within the fixed design domain $D$, two distinct level set functions $\phi_{pe}$ and $\phi_{sb}$ along with characteristic functions $\chi_{ps}$ and $\chi_{sp}$ signifying the respective domains are defined as follows:
$\left\{\begin{array}[]{ll}0<\phi_{pe}(\bm{x})\leq 1&\mathrm{if}~{}~{}\bm{x}\in\Omega_{pe}\\ \phi_{pe}(\bm{x})=0&\mathrm{if}~{}~{}\bm{x}\in\partial\Omega_{pe}\cup D_{sb}\\ -1\leq\phi_{pe}(\bm{x})<0&\mathrm{if}~{}~{}\bm{x}\in D_{pe}\setminus{\Omega}_{pe},\end{array}\right.$ (35)
$\left\{\begin{array}[]{ll}0<\phi_{sb}(\bm{x})\leq 1&\mathrm{if}~{}~{}\bm{x}\in\Omega_{sc}\\ \phi_{sb}(\bm{x})=0&\mathrm{if}~{}~{}\bm{x}\in\partial\Omega_{sc}\cup D_{pe}\\ -1\leq\phi_{sb}(\bm{x})<0&\mathrm{if}~{}~{}\bm{x}\in D_{sb}\setminus{\Omega}_{sb},\end{array}\right.$ (39)
$\left\{\begin{array}[]{ll}\chi_{ps}=1&\mathrm{in}~{}~{}D_{pe}\\ \chi_{ps}=0&\mathrm{in}~{}~{}D_{sb},\end{array}\right.$ (42)
$\left\{\begin{array}[]{ll}\chi_{sp}=0&\mathrm{in}~{}~{}D_{pe}\\ \chi_{sp}=1&\mathrm{in}~{}~{}D_{sb}.\end{array}\right.$ (45)
Each level set function for the piezoelectric and substrate domains is updated as follows:
$\frac{\partial\phi_{pe}}{\partial t}=-K(-\tilde{c}{F}^{\prime}_{pe}-\nabla(\bm{\tau}_{p}\nabla\phi_{pe})),$ (46)
$\frac{\partial\phi_{sb}}{\partial t}=-K(-\tilde{c}{F}^{\prime}_{sb}-\nabla(\bm{\tau}_{sb}\nabla\phi_{sb})),$ (47)
@@
where ${F}^{\prime}_{pe}$ is the gradient of the objective function $F_{pe}$ with respect to $\phi_{pe}$; ${F}^{\prime}_{sb}$ is the gradient of the objective function $F_{sb}$ with respect to $\phi_{sb}$; and $\bm{\tau}_{pe}$ and $\bm{\tau}_{sb}$ are the regularization factor tensors for $\phi_{pe}$ and $\phi_{sb}$, respectively. By individually defining the regularization coefficient tensors for the piezoelectric and substrate domains, the complexity in the thickness direction for each domain can be controlled, thereby enhancing the manufacturability.
@@ 3.2 Constraint: no piezoelectric material is placed without a substrate
Herein, we elucidate an approach that allows the imposition of geometric constraints using a fictitious physical function field denoted by $p$. The concept of fictitious physical models described by a steady-state anisotropic advection-diffusion equation was initially introduced by Sato et al. Sato et al. (2017). A further elaboration of the fictitious physical model using the anisotropic diffusion equation was provided by Yamada et al. Yamada and Noguchi (2022), which is described below.
@@
$\begin{split}-\text{div}(\bm{\kappa}_{p}\nabla p)+\chi_{sp}(p-\bar{p}_{s}(2\chi(\phi_{sb})-1))=0&\qquad\text{in}\quad D,\\ p=-\bar{p}_{0}&\qquad\text{on}\quad\Gamma_{p},\\ \nabla p\cdot\bm{n}=0&\qquad\text{on}\quad\partial D\backslash\Gamma_{p},\end{split}$ (48)
$\bm{\kappa}_{p}=\left[\begin{array}[]{lll}\kappa_{x}&0&0\\ 0&\kappa_{y}&0\\ 0&0&\kappa_{z}\end{array}\right],$ (52)
@@
where $\bm{\kappa}_{p}$ represents the conductivity tensor of $p$; $\kappa_{x}$, $\kappa_{y}$, and $\kappa_{z}$ represent the anisotropic conductivities along the $x$, $y$, and $z$ directions, respectively; and $\bar{p}_{s}\in\mathbb{R}_{+}$ and $\bar{p}_{0}\in\mathbb{R}_{+}$ are the constants for the generation and boundary conditions of $p$. In this study, the thickness direction was represented by the $z$-axis, which was employed as a constraint to determine the presence or absence of the substrate. The anisotropic conductivities $\kappa_{x}$ and $\kappa_{y}$ were set to values substantially lower than $\kappa_{z}$. The boundary $\Gamma_{p}$, subjected to the Dirichlet boundary condition, was identified as the boundary opposite to domain $D_{pe}$ within the domain $D_{sb}$. The second term in the initial equation of (48) ensures that $p=\bar{p}_{s}$ within the substrate material domain $\Omega_{sb}$. Within the substrate material domain $\Omega_{sb}$, the fictitious heat $p$ is generated and propagated in the $z$-direction towards the domain $D_{pe}$ when a material is present within $\Omega_{sb}$. Conversely, in the absence of a material in $\Omega_{sb}$, the boundary condition $p=-p_{0}$ is conveyed to the domain $D_{pe}$ owing to the lack of $p$ generation in $\Omega_{sb}$.
@@ 3.3 Formulation of the optimal design with proposed constraints
$\begin{split}\underset{\phi_{pe}}{\text{inf}}\qquad&F_{pe}(\omega_{oc},\omega_{sc})\\ \underset{\phi_{sb}}{\text{inf}}\qquad&F_{sb}(\omega_{oc},\omega_{sc})\\ \text{subject}\;\text{to}\qquad&a_{oc}(\phi_{pe},\phi_{sb},\omega_{oc},\bm{u}_{oc},\bm{v}_{oc},\varphi_{oc},p)=0,\\ &b_{oc}(\phi_{pe},\phi_{sb},\bm{u}_{oc},\varphi_{oc},v_{\varphi_{oc}},p)=0,\\ &a_{sc}(\phi_{pe},\phi_{sb},\omega_{sc},\bm{u}_{sc},\bm{v}_{sc},p)=0,\\ &a_{p}(\phi_{pe},\phi_{sb},p,v_{p})=0,\\ &G_{V}(\phi_{pe},\varphi_{oc})\leq 0,\end{split}$ (53)
@@
$\begin{split}&a_{oc}(\phi_{pe},\phi_{sb},\omega_{oc},\bm{u}_{oc},\bm{v}_{oc},\varphi_{oc})\\ &\quad=\biggl{(}\int_{\Omega_{pe}}\bm{s}(\bm{v}_{oc})^{T}\bm{C}^{E}_{pe}\bm{s}(\bm{u}_{oc})\ \text{d}\Omega-\omega_{oc}^{2}\int_{\Omega_{pe}}\rho_{pe}\bm{v}_{oc}^{T}\bm{u}_{oc}\ \text{d}\Omega\\ &\qquad+\int_{\Omega_{pe}}\bm{s}(\bm{v}_{oc})^{T}\bm{e}^{T}\bm{\nabla}\varphi_{oc}\ \text{d}\Omega\biggr{)}\\ &\qquad+\biggl{(}\int_{\Omega_{sb}}\bm{s}(\bm{v}_{oc})^{T}\bm{C}^{E}_{sb}\bm{s}(\bm{u}_{oc})\ \text{d}\Omega-\omega_{oc}^{2}\int_{\Omega_{sb}}\rho_{sb}\bm{v}_{oc}^{T}\bm{u}_{oc}\ \text{d}\Omega\biggr{)}=0,\end{split}$ (54)
@@
$\begin{split}&b_{oc}(\phi_{pe},\phi_{sb},\bm{u}_{oc},\varphi_{oc},v_{\varphi_{oc}})\\ &\quad=\int_{\Omega_{pe}}\bm{\nabla}v_{\varphi_{oc}}^{T}\bm{e}\bm{s}(\bm{u}_{oc})\ \text{d}\Omega-\int_{\Omega_{pe}}\bm{\nabla}v_{\varphi_{oc}}^{T}\bm{\varepsilon}^{S}\bm{\nabla}\varphi_{oc}\ \text{d}\Omega=0,\end{split}$ (55)
@@
$\begin{split}&a_{sc}(\phi_{pe},\phi_{sb},\omega_{sc},\bm{u}_{sc},\bm{v}_{sc})\\ &\quad=\biggl{(}\int_{\Omega_{pe}}\bm{s}(\bm{v}_{sc})^{T}\bm{C}^{E}_{pe}\bm{s}(\bm{u}_{sc})\ \text{d}\Omega-\omega_{sc}^{2}\int_{\Omega_{pe}}\rho_{pe}\bm{v}_{sc}^{T}\bm{u}_{sc}\ \text{d}\Omega\biggr{)}\\ &\qquad+\biggl{(}\int_{\Omega_{sb}}\bm{s}(\bm{v}_{sc})^{T}\bm{C}^{E}_{sb}\bm{s}(\bm{u}_{sc})\ \text{d}\Omega-\omega_{sc}^{2}\int_{\Omega_{sb}}\rho_{sb}\bm{v}_{sc}^{T}\bm{u}_{sc}\ \text{d}\Omega\biggr{)}=0,\end{split}$ (56)
$\begin{split}&a_{p}(\phi_{pe},\phi_{sb},p,v_{p})=\int_{D}\nabla v_{p}\bm{\kappa}_{p}\nabla p\ \text{d}\Omega+\int_{D}\chi_{sp}(p-\bar{p}_{s}(2\chi(\phi_{sb})-1))v_{p}\ \text{d}\Omega=0,\end{split}$ (57)
@@
$\begin{split}&G_{V}(\phi_{pe},\varphi)=\int_{\Omega_{pe}}\ \text{d}\Omega-\frac{\varepsilon_{z}L_{z}^{2}\int_{\Omega_{pe}}\bm{n}_{z}\cdot\nabla\varphi\ \text{d}\Omega}{\bar{V}_{minV}}\leq 0,\end{split}$ (58)
@@
where $\bm{C}^{E}_{pe}$ and $\bm{C}^{E}_{sb}$ are the elastic tensors of the piezoelectric and substrate materials, respectively. Similarly, $\rho_{pe}$ and $\rho_{sb}$ represent the densities of the piezoelectric and substrate materials, respectively. Additionally, $v_{p}$ acts as a test function for the fictitious physical function $p$. The potential, symbolized by $\varphi$, is obtained from the displacement $\bm{u}$ derived via the modal superposition method. This is distinct from $\varphi_{oc}$, which is determined based on the eigenvector $\bm{u}_{oc}$. A comprehensive exposition on the derivation of the potential $\varphi$ through the modal superposition method can be found in A.

### v2

@@ 3.1 Constraint: the same cross-sectional shape
3.1.1 Constraint: the same cross-sectional shape throughout the design domain
@@
Within the framework of topology optimization obtained via the level set method, qualitative control over complexity can be regulated by the regularization coefficient $\tau$, as defined in Eq. (39) [32, 56]. To leverage this control over complexity, we adopt the approach used by Yamada et al. [32] to assign regularization coefficients that curtail the complexity in the thickness direction of the structure, thereby facilitating manufacturability via etching. In this study, we extend the regularization coefficient to a tensor $\bm{\tau}$ and modify Eqs. (39) and (41) as follows:
$\begin{split}{\underset{\phi}{\text{inf}}}\qquad&{F_{R}(u,\phi)=\int_{D}f(u)\chi_{\phi}(\phi)\mathrm{d\Omega}+\int_{D}\frac{1}{2}|(\nabla\phi)^{T}\bm{\tau}\nabla\phi|\ \text{d}\Omega}\\ \text{subject to}\qquad&\text{the governing equations of the physical domain},\\ &\text{additional constraint equations,}\end{split}$ (42)
@@
To maintain a consistent cross-sectional shape throughout the design domain, the optimization problem is formulated using a single level set function $\phi$ and a single weight factor $\alpha$. In this study, since the physical properties of the object domain differ between the design domain of the piezoelectric material $D_{pe}$ and the design domain of the substrate $D_{sb}$, a constant function $\phi_{ps}$ is used to define the governing equations as follows:
$\left\{\begin{array}[]{ll}\phi_{ps}(\bm{x})=1&\mathrm{if}~{}~{}\bm{x}\in D_{pe}\\ \phi_{ps}(\bm{x})=-1&\mathrm{if}~{}~{}\bm{x}\in D_{sb},\end{array}\right.$ (50)
$\begin{array}[]{rll}\underset{\phi}{\text{inf}}&F=\alpha F_{k}+(1-\alpha)F_{\omega},&\\ \text{subject to}&{(\bm{C}^{E}_{pe}\chi_{\phi}(\phi_{ps})+\bm{C}^{E}_{sb}\chi_{\phi}(1-\phi_{ps}))\chi_{\phi}(\phi)\bm{s}(\bm{u}_{oci})}&\\ &{\quad-\omega_{oci}^{2}(\rho_{pe}\chi_{\phi}(\phi_{ps})+\rho_{sb}(1-\chi_{\phi}(\phi_{ps})))\chi_{\phi}(\phi)\bm{u}_{oci}}&\\ &{\quad+\bm{e}^{T}\chi_{\phi}(\phi_{ps})\bm{\nabla}\varphi_{oci}=0}&\quad{\text{in}\;D,}\\[5.0pt] &{\bm{e}\chi_{\phi}(\phi_{ps})\chi_{\phi}(\phi)\bm{s}(\bm{u}_{oci})-\bm{\varepsilon}^{S}\chi_{\phi}(\phi)\bm{\nabla}\varphi_{oci}=0}&\quad{\text{in}\;D,}\\[5.0pt] &{(\bm{C}^{E}_{pe}\chi_{\phi}(\phi_{ps})+\bm{C}^{E}_{sb}\chi_{\phi}(1-\phi_{ps}))\chi_{\phi}(\phi)\bm{s}(\bm{u}_{sci})}&\\ &{\quad-\omega_{sci}^{2}(\rho_{pe}\chi_{\phi}(\phi_{ps})+\rho_{sb}(1-\chi_{\phi}(\phi_{ps})))\chi_{\phi}(\phi)\bm{u}_{sci}=0}&\quad{\text{in}\;D.}\end{array}$ (58)
3.1.2 Constraint: the same cross-sectional shape in each domain
@@
Although the method described above controls the complexity of the structure throughout the design domain, the substrate and piezoelectric components can be readily processed via etching, even if their structures differ. It is desirable to control the complexity in the thickness direction within the substrate and piezoelectric domains without affecting the structural differences at the interface of each domain. To this end, in this study, we defined separate level set functions for the substrate and piezoelectric structures, invoking the extended level set method [35], which is capable of handling multiple materials and updating each level set function concurrently. In the extended level set method, when dealing with a void domain and two types of object domains, making three types of domains in total, the level set function matrix is defined as follows:
$\bm{\phi}=\left[\begin{array}[]{lll}\phi_{00}&\phi_{01}&\phi_{02}\\ \phi_{10}&\phi_{11}&\phi_{12}\\ \phi_{20}&\phi_{21}&\phi_{22}\end{array}\right],$ (62)
$\chi_{a}$ $=\prod_{b\in[0,1,2]\backslash a}\chi_{\phi}(\phi_{ba}),$ (63)
$\phi_{aa}$ $=0\qquad(a\in[0,1,2])$ (64)
$\phi_{ab}$ $=-\phi_{ba}\qquad(a\neq b)$ (65)
In this framework, $\phi_{p\psi}$ is defined as $\phi_{10}$, representing the boundary between the piezoelectric and void domains, and $\phi_{s\psi}$ is defined as $\phi_{20}$, representing the boundary between the substrate and void domains. Therefore, the level set function matrix can be redefined as follows:
$\bm{\phi}$ $=\left[\begin{array}[]{lll}\phi_{\psi\psi}&\phi_{\psi p}&\phi_{\psi s}\\ \phi_{p\psi}&\phi_{pp}&\phi_{ps}\\ \phi_{s\psi}&\phi_{sp}&\phi_{ss}\end{array}\right],$ (69)
$\phi_{\psi p}$ $=-\phi_{p\psi},$ (70)
$\phi_{\psi s}$ $=-\phi_{s_{\psi}},$ (71)
$\phi_{sp}$ $=-\phi_{ps},$ (72)
$\phi_{\psi\psi}$ $=\phi_{pp}=\phi_{ss}=0,$ (73)
$\chi_{p}(\bm{x})$ =$=\left\{\begin{array}[]{ll}1&\mathrm{if}~{}~{}\bm{x}\in\Omega_{pe}\\ 0&\mathrm{if}~{}~{}\bm{x}\in D\setminus{\Omega}_{pe},\end{array}\right.$ (76)
$\chi_{s}(\bm{x})$ =$=\left\{\begin{array}[]{ll}1&\mathrm{if}~{}~{}\bm{x}\in\Omega_{sb}\\ 0&\mathrm{if}~{}~{}\bm{x}\in D\setminus{\Omega}_{sb},\end{array}\right.$ (79)
$\chi_{\psi}(\bm{x})$ =$=\left\{\begin{array}[]{ll}1&\mathrm{if}~{}~{}\bm{x}\in D\setminus(\Omega_{pe}\cup\Omega_{sb})\\ 0&\mathrm{if}~{}~{}\bm{x}\in\Omega_{pe}\cup\Omega_{sb},\end{array}\right.$ (82)
where $\chi_{p}$ denotes the piezoelectric material domain, $\chi_{s}$ denotes the substrate material domain, and $\chi_{\psi}$ denotes void domain. The fixed design domains for the substrate and piezoelectric material, denoted as $D_{sb}$ and $D_{pe}$, respectively, are assumed to remain invariant throughout the optimization procedure. The entire fixed design domain, denoted as $D$, encompasses both $D_{pe}$ and $D_{sb}$. The level set function $\phi_{ps}$, representing the boundary between the domains $\omega_{pe}$ and $\omega_{sb}$, is defined as a constant value during the optimization process, as shown in the following equation:
$\left\{\begin{array}[]{ll}\phi_{ps}(\bm{x})=1&\mathrm{if}~{}~{}\bm{x}\in D_{pe}\\ \phi_{ps}(\bm{x})=-1&\mathrm{if}~{}~{}\bm{x}\in D_{sb},\end{array}\right.$ (85)
The level set functions $\phi_{p\psi}$ and $\phi_{s\psi}$ are updated as follows:
$\frac{\partial\phi_{p\psi}}{\partial t}=-K\left(-\tilde{c}F^{\prime}_{p\psi}-\nabla(\bm{\tau}_{p\psi}\nabla\phi_{p\psi})\right),$ (86)
$\frac{\partial\phi_{s\psi}}{\partial t}=-K\left(-\tilde{c}F^{\prime}_{s\psi}-\nabla(\bm{\tau}_{s\psi}\nabla\phi_{s\psi})\right),$ (87)
@@
where $F^{\prime}_{p\psi}$ and $F^{\prime}_{s\psi}$ are the gradients of the objective function with respect to $\phi_{p\psi}$ and $\phi_{s\psi}$, respectively, and $\bm{\tau}_{p\psi}$ and $\bm{\tau}_{s\psi}$ are the regularization coefficient tensors for $\phi_{p\psi}$ and $\phi_{s\psi}$, respectively. By individually defining the regularization coefficient tensors for the piezoelectric and substrate domains, the complexity in the thickness direction for each domain can be controlled. This approach allows for an increase in design flexibility without compromising manufacturability, compared to using a single level set function for the entire fixed design domain $D$.
@@ 3.2 Constraint: no piezoelectric material is placed without a substrate
Herein, we elucidate an approach that allows the imposition of geometric constraints by using a fictitious physical function field denoted by ${\xi}$. The concept of fictitious physical models described by a steady-state anisotropic advection-diffusion equation is initially introduced by Sato et al. [28]. A further elaboration of the fictitious physical model using the anisotropic diffusion equation is provided by Yamada et al. [30], which is described below.
@@
$\begin{split}{-\text{div}(\bm{\kappa}_{{\xi}}\nabla{\xi})=0}&{\qquad\text{in}\quad D,}\\ {{\xi}=\bar{{\xi}}_{s}}&{\qquad\text{in}\quad\Omega_{sb},}\\ {\xi}=-\bar{{\xi}}_{0}&\qquad\text{on}\quad\Gamma_{\xi},\\ \nabla{\xi}\cdot\bm{n}=0&\qquad\text{on}\quad\partial D\backslash\Gamma_{\xi},\end{split}$ (88)
$\bm{\kappa}_{\xi}=\left[\begin{array}[]{lll}\kappa_{x}&0&0\\ 0&\kappa_{y}&0\\ 0&0&\kappa_{z}\end{array}\right],$ (92)
@@
where $\partial D$ denotes the outer boundary of the entire fixed design domain $D$; $\bm{\kappa}_{\xi}$ represents the conductivity tensor of ${\xi}$; $\kappa_{x}$, $\kappa_{y}$, and $\kappa_{z}$ represent the anisotropic conductivities along the $x$, $y$, and $z$ directions, respectively; and $\bar{{\xi}}_{s}\in\mathbb{R}_{+}$ and $\bar{{\xi}}_{0}\in\mathbb{R}_{+}$ are the constants for the domain and boundary conditions of ${\xi}$. In this study, the thickness direction is represented by the $z$-axis, which is employed as a constraint to determine the presence or absence of the substrate. The anisotropic conductivities $\kappa_{x}$ and $\kappa_{y}$ are set to values substantially lower than $\kappa_{z}$. The boundary $\Gamma_{\xi}$, subjected to the Dirichlet boundary condition, is identified as the boundary opposite to domain $D_{pe}$ within the domain $D_{sb}$. The second equation of (88) ensures that ${\xi}=\bar{{\xi}}_{s}$ within the substrate material domain $\Omega_{sb}$. The physical field ${\xi}$ behaves as a fictitious heat within the fixed design domain $D$. In the domain $\Omega_{sb}$, it acts as a heat source with a temperature ${\xi}=\bar{{\xi}}_{s}$, and a cold heat bath with ${\xi}=-\bar{{\xi}}_{0}$ is applied at the boundary $\Gamma_{\xi}$. Within the substrate material domain $\Omega_{sb}$, the fictitious heat ${\xi}$ is applied and propagates in the $z$-direction towards the domain $D_{pe}$ when a material is present within $\Omega_{sb}$. Conversely, in the absence of a material within $D_{sb}$, the boundary condition ${\xi}=-\bar{\xi}_{0}$ is conveyed to the domain $D_{pe}$ owing to the lack of ${\xi}$ propagation in $\Omega_{sb}$. Consequently, within the piezoelectric design domain $D_{pe}$, ${\xi}$ assumes positive values above the substrate and negative values elsewhere. By defining $\Omega_{pe}$ as regions where $\phi_{p\psi}>0$ and ${\xi}>0$, we ensure that the structure adheres to the substrate-dependent constraint:
$\chi_{p}(\bm{x})=\left\{\begin{array}[]{ll}1\qquad\mathrm{for}\quad\bm{x}\in\Omega_{pe}\\ 0\qquad\mathrm{for}\quad\bm{x}\in D\backslash\Omega_{pe},\\ \end{array}\right.$ (95)
$\chi_{p}(\bm{x}):=\chi_{\phi}(\phi_{p\psi}(\bm{x}))\chi_{\phi}(\xi(\bm{x}))$ (96)
Figure 6 visually elucidates the behavior of the fictitious heat field ${\xi}$ and its role in defining the piezoelectric material domain. The figure depicts the initial structure without constraints, followed by the propagation of the fictitious heat field within both the substrate and piezoelectric domains. The fictitious heat field ensures that no piezoelectric material is placed without the presence of a substrate.
Figure 6: Schematic illustration of the fictitious heat field used to enforce the substrate-dependent constraint. Non-compliant Structure: Initial configuration without constraints. Fictitious Heat Bath at $\Gamma_{\xi}$: Cold heat bath with $\xi=-\bar{\xi}_{0}$ is applied at $\Gamma_{\xi}$, the boundary opposite to domain $D_{pe}$ within domain $D_{sb}$. Fictitious Heat Field in $D_{sb}$: Substrate material domain $\Omega_{sb}$ as a heat source with $\xi=\bar{\xi}_{s}$. Fictitious Heat Propagation: Heat propagation in the $z$-direction towards $D_{pe}$. In the absence of material in $\Omega_{sb}$, $\xi=-\bar{\xi}_{0}$ is conveyed to $D_{pe}$. Piezoelectric Domain Represented by Fictitious Heat Field: $\Omega_{pe}$ expressed as $\chi_{p}(\bm{x}):=\chi_{\phi}(\phi_{p\psi}(\bm{x}))\chi_{\phi}(\xi(\bm{x}))$ to satisfy the substrate-dependent constraint. Compliant Structure: Structure satisfying the substrate-dependent constraint.
@@ 3.3 Formulation of the optimal design with proposed constraints
$\begin{split}\underset{{\phi_{p\psi}}}{\text{inf}}\qquad&F_{pe}\\ \underset{{\phi_{s\psi}}}{\text{inf}}\qquad&F_{sb}\\ \text{subject}\;\text{to}\qquad&\text{for}\;i=1,\ldots,n\\ &a_{oci}({\phi_{p\psi}},{\phi_{s\psi}},\omega_{oci},\bm{u}_{oci},\bm{v}_{oci},\varphi_{oci},{\xi})=0,\\ &b_{oci}({\phi_{p\psi}},{\phi_{s\psi}},\bm{u}_{oci},\varphi_{oci},v_{\varphi_{oci}},{\xi})=0,\\ &a_{sci}({\phi_{p\psi}},{\phi_{s\psi}},\omega_{sci},\bm{u}_{sci},\bm{v}_{sci},\xi)=0,\\ &a_{\xi}({\phi_{p\psi}},{\phi_{s\psi}},{\xi},v_{\xi})=0,\\ &G_{V}({\phi_{p\psi}},\varphi_{n})\leq 0,\end{split}$ (97)
@@
$\begin{split}&a_{oci}({\phi_{p\psi}},{\phi_{s\psi}},\omega_{oci},\bm{u}_{oci},\bm{v}_{oci},\varphi_{oci})\\ &\quad=\biggl{(}\int_{\Omega_{pe}}\bm{s}(\bm{v}_{oci})^{T}\bm{C}^{E}_{pe}\bm{s}(\bm{u}_{oci})\ \text{d}\Omega-\omega_{oc}^{2}\int_{\Omega_{pe}}\rho_{pe}\bm{v}_{oci}^{T}\bm{u}_{oci}\ \text{d}\Omega\\ &\qquad+\int_{\Omega_{pe}}\bm{s}(\bm{v}_{oci})^{T}\bm{e}^{T}\bm{\nabla}\varphi_{oci}\ \text{d}\Omega\biggr{)}\\ &\qquad+\biggl{(}\int_{\Omega_{sb}}\bm{s}(\bm{v}_{oci})^{T}\bm{C}^{E}_{sb}\bm{s}(\bm{u}_{oci})\ \text{d}\Omega-\omega_{oc}^{2}\int_{\Omega_{sb}}\rho_{sb}\bm{v}_{oci}^{T}\bm{u}_{oci}\ \text{d}\Omega\biggr{)}=0,\end{split}$ (98)
@@
$\begin{split}&b_{oci}({\phi_{p\psi}},{\phi_{s\psi}},\bm{u}_{oci},\varphi_{oci},v_{\varphi_{oci}})\\ &\quad=\int_{\Omega_{pe}}\bm{\nabla}v_{\varphi_{oci}}^{T}\bm{e}\bm{s}(\bm{u}_{oci})\ \text{d}\Omega-\int_{\Omega_{pe}}\bm{\nabla}v_{\varphi_{oci}}^{T}\bm{\varepsilon}^{S}\bm{\nabla}\varphi_{oci}\ \text{d}\Omega=0,\end{split}$ (99)
@@
$\begin{split}&a_{sci}({\phi_{p\psi}},{\phi_{s\psi}},\omega_{sci},\bm{u}_{sci},\bm{v}_{sci})\\ &\quad=\biggl{(}\int_{\Omega_{pe}}\bm{s}(\bm{v}_{sci})^{T}\bm{C}^{E}_{pe}\bm{s}(\bm{u}_{sci})\ \text{d}\Omega-\omega_{sci}^{2}\int_{\Omega_{pe}}\rho_{pe}\bm{v}_{sci}^{T}\bm{u}_{sci}\ \text{d}\Omega\biggr{)}\\ &\qquad+\biggl{(}\int_{\Omega_{sb}}\bm{s}(\bm{v}_{sci})^{T}\bm{C}^{E}_{sb}\bm{s}(\bm{u}_{sci})\ \text{d}\Omega-\omega_{sci}^{2}\int_{\Omega_{sb}}\rho_{sb}\bm{v}_{sci}^{T}\bm{u}_{sci}\ \text{d}\Omega\biggr{)}=0,\end{split}$ (100)
$\begin{split}&a_{\xi}({\phi_{p\psi}},{\phi_{s\psi}},{\xi},v_{\xi})=\int_{D}\nabla v_{\xi}\bm{\kappa}_{\xi}\nabla{\xi}\ \text{d}\Omega+\int_{D}{h(\phi_{sp})}({\xi}-\bar{{\xi}}_{s}(2\chi({\phi_{s\psi}})-1))v_{\xi}\ \text{d}\Omega=0,\end{split}$ (101)
@@
$\begin{split}&G_{V}({\phi_{p\psi}},\varphi_{n})=\int_{\Omega_{pe}}\ \text{d}\Omega-\frac{\varepsilon_{z}L_{z}^{2}\int_{\Omega_{pe}}\bm{n}_{z}\cdot\nabla\varphi_{n}\ \text{d}\Omega}{\bar{V}_{minV}}\leq 0,\end{split}$ (102)
@@
where $\rho_{pe}$ and $\rho_{sb}$ represent the densities of the piezoelectric and substrate materials, respectively, whereas $\bm{v}_{oci}$, $\bm{v}_{sci}$, $v_{\varphi oci}$, and $v_{\xi}$ serve as test functions. The test functions are detailed as follows: $\bm{v}_{oci}$ and $\bm{v}_{sci}$ are test functions for the eigenvectors $\bm{u}_{oci}$ and $\bm{u}_{sci}$, respectively. $v_{\varphi oci}$ is a test function for $\varphi_{oci}$. $v_{\xi}$ is a test function for the fictitious physical function $\xi$. The potential, symbolized by $\varphi_{n}$, is obtained from the displacement $\bm{u}$ derived via the modal superposition method. This is distinct from $\varphi_{oci}$, which is determined based on the eigenvector $\bm{u}_{oci}$. A comprehensive exposition on the derivation of the potential $\varphi_{n}$ through the modal superposition method can be found in A.